MARS DESERT RESEARCH STATION

Sol Summary – April 20th

Crew 178 Daily Summary Report 20APR2017
 MDRS Daily Summary Report for sol 10
Summary Title: 11th EVA
Mission Status:  Good

Sol Activity Summary:

Hab
08.00:   Breakfast
13.00: Lunch
15.00: Muscle wasting exercises experiment

GreenHab
/

ScienceDome
Experiment in biology: survey of plants growth and health progress
Experiment reports redaction

EVA
09.00: UTM NAD27 12S 4255500-518500
Diffractometer study: sample collection
Telecommunication relays experiment
Testing and installing sound beacons

Weather:  Sunny

Crew Physical Status: Everyone is fine

Look Ahead Plan: Last EVA and concluding experiments

Reports to be file: Operation Report, Daily Summary Report, EVA 11 Report, Journalist Report

Support Requested:
/

Sol Summary – April 18th

MDRS Daily Summary Report for sol 8

Summary Title: 9th EVA

Mission Status:  Good

Sol Activity Summary:

Hab
8.00:   Breakfast
10.00: Muscle wasting exercises
13.00: Lunch

GreenHab
/

ScienceDome
Experiment reports redaction

EVA
10.30: UTM NAD27 12S 4251500-518500
Telecommunication relays test
Diffractometer study: data collection

Weather:
Sunny and strong wind

Crew Physical Status
Everyone is fine

Look Ahead Plan:
EVA + Experiments

Reports to be file: Operation Report, Daily Summary Report, Commander Report, EVA 9 Report

Support Requested:

Sol Summary – April 17th

Crew 178 Daily Summary Report 17APR2017

 

MDRS Daily Summary Report for sol 7

 

Summary Title: 7th and 8th EVA

 

Mission Status:  Good

 

Sol Activity Summary:

 

Hab

8.00:   Breakfast

13.00: Lunch

14.30: Preparation of the last telecommunication relays’s EVA

 

GreenHab

/

 

ScienceDome

Experiment reports redaction

Muons experiment: tests on the scintillator

 

EVA

9.00 UTM NAD27 12S 4255500-518500

X-ray diffraction study: collection of data

Positioning of sound beacons for mapping of dangerous area

Topography mapping

 

14.00: UTM NAD27 12S 4251500-518500

Topography mapping

Issue with the probe

 

Weather:

Sunny

 

Crew Physical Status

Everyone is fine

 

Look Ahead Plan:

EVA + Experiments

 

 

Reports to be file: Operation Report, Daily Summary Report, Journalist Report, HSO Report

 

Support Requested: /

Sol Summary – April 16th

Crew 178 Daily Summary Report 16APR2017

MDRS Daily Summary Report for sol 6

Summary Title: Sixth EVA

Mission Status:  Good

Sol Activity Summary:

Hab
8.30:   Breakfast
13.00: Lunch
16.00: Muscle wasting experiment: cardio exercise session
18:30: Daily group meeting
19:00: CapCom
20:30: Dinner

GreenHab
/

ScienceDome
/

EVA
10.30 UTM NAD27 12S 4255500-518500
Media: pictures and videos of north area
Telecommunication relays experiment

Weather:
Sunny

Crew Physical Status
Everyone is fine

Look Ahead Plan:
EVA + Experiments

Reports to be file: Operations Report, Daily Summary Report, Commander Report, EVA #6 report

Support Requested:
/

Sol Summary – April 15th

Crew 178 Daily Summary Report 15APR2017

 

MDRS Daily Summary Report for sol 5

 

Summary Title: Fifth EVA

 

Mission Status:  Good

 

Sol Activity Summary:

 

Hab

8.30:   Breakfast

11.00: Muscle wasting experiment: exercise session

12.30: Lunch

 

GreenHab

/

 

ScienceDome

14.00: End setup recycling experiment. 

 

EVA

14.00 UTM NAD27 12S 4247631-519930

Ground penetrating radar experiment : data collection

Topography Geomatic study: fixing probe

Mapping of dangerous areas with the use of probes

 

Weather:

Sunny

 

Crew Physical Status:

Everyone is fine

 

Look Ahead Plan:

EVA + Experiments

 

 

Reports to be file: Operation Report , Daily Summary Report, HSO Report, Journalist Report

 

Support Requested:

Information about water consumption

Bottom part of the hab outside repair

What sould we do with the inside toilet ?

Sol Summary – April 14th

Crew 178 Daily Summary Report 14APR2017

 

MDRS Daily Summary Report for sol 4

 

Summary Title: Setups for several experiments and fourth EVA

 

Mission Status:  Good

 

Sol Activity Summary:

 

Hab

8.30 : Breakfast at 8.30

10.00: Muscle wasting experiment: exercise session

12.30: Lunch

 

GreenHab

9.30: Watering rice plants for recycling experiment

 

ScienceDome

10.30: Setup recycling experiment : hydroponic solution with different urea concentrations and plants setting in the incubator.  

 

EVA

16.30 UTM NAD27 12S 4251500/518500

Ground penetrating radar experiment : data collection

Topography Geomatic study : data collection

Omnidirectional telecommunication relays : installation of a first relay

Media: pictures and videos of the different activities

Issues : ATV (out of gas) and faulty probe

 

Weather:

Sunny and slightly windy

 

Crew Physical Status

Everyone is fine

 

Look Ahead Plan:

EVA + Experiments

 

 

Reports to be file: Operation Report , Daily Summary Report, HSO Report, Astronomy Report, Journalist Report

 

Support Requested:

Musk Observatory : procedure to align mirrors

Sol Summary – April 13th

Crew 178 Daily Summary Report 13 APR 2017

MDRS Daily Summary Report for sol 3

Prepared by: Aurian d’Avernas, Commander, Crew 178

Summary Title: Third EVA, fixing water issues and preparations of experiments

Mission Status:  Good

Sol Activity Summary: See Commander Report

Look Ahead Plan: We will perform an EVA for 4 experiments and try to fix restroom and water heater issues.

Anomalies in work:  Inside restroom present problems, while the outside one fell down due to heavy gusts of wind.

Weather: Sunny and windy

Crew Physical Status: Good

EVA: In the North UTM NAD27 425300 – 518000

Reports to be file: Commander Report + Daily Summary Report + Journalist Report + Engineer Report + EVA report + EVA request

Support Requested: Need assistance for restrooms and the water heater issues (see Engineer report)

Sol Summary – April 12th

Crew 178 Daily Summary Report 12 APR 2017
MDRS Daily Summary Report for sol 2
Summary Title: Second EVA, preparations for other experiments and Internet fixing
Mission Status: Good
Sol Activity Summary:
“Today, we again woke up at 8 am after having enjoyed “The Martian” movie yesterday night with the whole team. In the morning, Nathalie invented the “Pancake dance” while she was cooking delicious pancakes for everyone, although her first 5 pancakes were closer to sharcoal than cake. Elke finished the setup of the last parameters for her second EVA using her ground penetrating radar. Calogero continued to configure his software applications for his geomatics EVA. At noon, we ate strange canned turkey meat and a Minestrone bear creek soup with our favorite bread, oven-cooked with love by our Crew Scientist, Mathieu. On the whole, this meal was still a pleasure. In the afternoon, the second EVA was launched under a very sunny weather reaching about 26 degrees Celsius. A fun fact about our Biohazard space suits is that they’re cloaked with an external plastic layer, allowing us to sweat like never inside. Extreme conditions are fulfilled. Quentin and Elke will try to reconstruct a Shepherd’s pie with lyophilized food for an amazing dinner tonight. After dinner, every crew member will have still hard work to finish preparing at best the next days.
Aurian d’Avernas, the Crew Commander”
Look Ahead Plan: The HughesNet technician came today to fix the parable with concrete. We identified that some laptops were automatically burning our daily Internet data with the automatic Windows updates which are very tricky to disable. We finally found the solution and tomorrow will be the first day without any Internet issue. The preparations of the crew experiments are going well. Moreover, we will send our biographies and crew photos tomorrow morning without exception.
Anomalies in work: The wheel of the ground penetrating radar does not work properly. It should be fixed by a phone call to the company of this material tomorrow morning by our Crew Geologist.
Weather: Clear sky and very sunny.
Crew Physical Status: Very good.
EVA: 1 km at the North of the Hab.
Reports to be file: EVA #2 report + Journalist report + Engineer report + Daily Summary Report
Support Requested: none

Sol Summary – April 11th

Summary Title: Technical setups for several experiments and first EVA

Mission Status:  Good

Sol Activity Summary:

We woke up today at 8 am. The Crew Astronomer and the Crew Commander broke the sim for two hours (10-12 am) in order to take the aerial delivery at Hollow Mountain (5 days of late) containing our scientific equipment. We shared meal at 1 pm (tomato soup with bread). The Crew Geologist and the Crew Scientist started their EVA at 2 pm and came back at 3.30 pm. In the meanwhile, the other crew members set up their experiment in the Hab and the Science Lab. The Crew Geologist prepared its plants launched the GreenHab activities. We had to do an emergency EVA (in spacesuits of course) to the Shannon’s trailer from 7 to 9 pm in order to start CapCom session. Indeed, some electronic devices probably consumed all the daily Internet data for this day, that is why we were constrained to exceptionally go to the Shannon’s trailer. At 6.30 pm, we started a group meeting to debrief the day and prepare the day of tomorrow, and also to disable all the electronic devices of each crew member to avoid any further problem with the daily Internet datas. After the evening meal, some of the crew members enjoyed a movie, while others continued to work.Look Ahead Plan: (How are you going to follow up?  What’s the plan for tomorrow?)

Anomalies in work: Internet data were consumed par mégarde

Weather: The beginning of the day started to be dégagé and clear. In the afternoon, some clouds added and the wind became more violent ???

Crew Physical Status: Everyone is fine

EVA: (where did you go?)

Reports to be file: Crew Engineer Report, Journalist Report,

Support Requested: The tunnel structure between the Hab and the Science Lab has suffered from heavy gusts of wind, it will be difficult

Sol Summary – April 10th

Crew 178 Daily Summary Report  10/04/2017

MDRS Daily Summary Report for sol 0

Summary Title: Fixing Internet and aerial delivery delays

Mission Status:  Several technical problems prevent us to start the simulation

Sol Activity Summary:
The Crew Engineer and the Crew Commander started in the morning to fix the Hab’s Internet Connection: it has been a success since we identified that the issue was due to the rotation in the vertical axis of the parable. This rotation varied of a few degrees because of lots of gusts of wind. We then fixed this issue by adding a fastening system with ropes. The rest of the day was dedicated to discover the surrounding areas with ATV’s and rovers to have a better idea of the relief. We also started physical tests (known as the “bip test”) and other physical exercises outside in the framework of the experiment of our physiotherapist. Finally, we passed few hours to fix the aerial delivery delays by calling several customer services. We got the confirmation that our packages will be finally delivered tomorrow at 10 am.

Look Ahead Plan: Internet connection has been fixed and the delivery of our packages should coming tomorrow, we then plan to start the simulation tonight at 6 pm although two of our crew members will have to break the sim tomorrow morning to get our aerial delivery.

Anomalies in work:  The technical assistance of HughesNet should still come to install a better fastening system for the parable since our solution is not robust against strong winds.

Weather: The weather was cleared and the sun predominant.

Crew Physical Status: Perfect health

EVA: No EVA

Reports to be file: Crew Engineer Report + Daily Summary

Support Requested: